Write the dual of the following LPP after reducing it to canonical form.

Min Z = 3x1 + 4x2 + 3x3

Subject to

2x1+4x2 =12

5x1+3x3 ≥11

6x1+ x2 ≥ 8

x1,x2,x3≥0


1
Expert's answer
2021-06-17T11:44:05-0400

The LPP is given as;

3x1 + 4x2 + 3x3 = Min Z

2x1 + 4x2 + 0x3 =12

5x1 + 0x2 + 3x3 ≥ 11

6x1+ x2 + 0x3 ≥ 8

x1 ≥ 0

x2 ≥ 0

x3 ≥ 0


We construct the matrix of coefficients for the free problem, direct problem, and forward problem and name them A, B, C respectively, where;

A = "\begin{pmatrix}
 3 & 4 &3\\
 
\end{pmatrix}"

B= "\begin{pmatrix}
 12 \\
 
 11\\
8
\end{pmatrix}"

D ="\begin{pmatrix}
 2& 4&0\\
 5 & 0&3\\
6&1&0

\end{pmatrix}"


To find the dual of A, we compute the transpose matrix of B and to compute the dual of B, we find the tranpose matrix of A. The dual of C is the transpose matrix of C

A dual = "\begin{pmatrix}
 12 & 11 &8\\
 
\end{pmatrix}"

B dual = "\begin{pmatrix}
 3 \\
 
 4\\
3
\end{pmatrix}"

C dual = "\begin{pmatrix}
 2& 5&6\\
 4& 0&1\\
0&3&0

\end{pmatrix}"


To find dual of the problem, we maximize the objective function. Also, we have the same number of constraints in the dual problem as that of the direct problem.

A maximized objective function will contain either the inequality sign ""\le" " or " ="


Therefore, the dual LPP is given as follows;

12y1 +11y2 + 8y3 = Max Z

2y1 +5y2 + 6y3 "\le" 3

4y1 + y3 "\le" 4

3y2 "\le" 3


y1, y2, y3 ≥ 0